# Global-Decision-Focused Neural ODEs for Proactive Grid Resilience Management

### v1

@@ I Introduction
To proactively mitigate power grid disruptions, it is crucial to anticipate the long-term impact of extreme events on power systems and implement preemptive measures accordingly. Accurate forecasting enables utility companies and policymakers to assess risks, prioritize interventions, optimize grid reinforcement, and allocate resources efficiently. By identifying vulnerable regions in advance, decision-makers can develop strategies that not only reduce power disruptions but also strengthen overall grid resilience.
@@ I-A Related Work
Two critical problems in this domain illustrate the limitations of such approach. The generator deployment problem determines optimal DG placement and sizing to mitigate outages caused by extreme events [17, 18, 16]. Traditional approaches emphasize on optimization techniques like genetic algorithms and particle swarm optimization to minimize power losses and improve voltage stability [18]. Similarly, the power line undergrounding problem aims to reinforce the grid infrastructure by undergrounding transmission lines to withstand hazards such as hurricanes and earthquakes [23, 24]. Conventional methods prioritize static risk assessments using historical data, lacking adaptability to evolving threats [25, 26]. While these approaches improve resilience, they fail to account for dynamic outage scenarios, limiting their effectiveness in extreme events.
To address these limitations, this paper proposes integrating decision-focused learning (DFL) with predictive modeling. By embedding decision-making objectives directly into the learning process, our approach aligns predictions with optimization goals, improving resource allocation and grid reinforcement strategies. This integrated approach enables more adaptive and proactive decision-making, enhancing grid resilience against escalating natural-hazard-related risks.
Power Outage Modeling. Accurate power outage forecasting is crucial for enhancing grid reliability and resilience [27, 28]. Various machine learning and statistical methods have been employed to predict outages under different conditions. For example, some studies have focused on forecasting outage durations using historical natural language processing (NLP) data [29]. These approaches incorporate neural networks enriched with environmental factors and semantic analysis of field reports, providing real-time updates and enhancing predictive performance through text analysis.
@@ III-A Neural ODEs for Power Outages
Drawing inspiration from the Susceptible Infectious Recovered (SIR) models commonly used in epidemic modeling [50], we conceptualize power outages within a unit as the spread of a “virus”. In this analogy, outages propagate among customers due to local transmission line failures, while restorations provide lasting resistance to subsequent outages. We formalize this analogy with the following two assumptions: ($i$) The number of unaffected customers $U_{k}(t)$ decreases over time as some customers transition from being unaffected to experiencing outages. This transition is governed by the failure transmission rate, denoted by $\phi(\bm{z}_{k};\theta_{U})$, which represents the probability of an outage per contact multiplied by the rate of such contacts, depending on its connectivity in the grid network. ($ii$) Conversely, the number of customers with restored power ($R_{k}(t)$) gradually increases as the system operator repairs transmission lines and restores service. This process is driven by the restoration rate, denoted $\phi(\bm{z}_{k};\theta_{R})$. Both transmission and restoration rates, $\phi(\bm{z}_{k};\theta_{U})$ and $\phi(\bm{z}_{k};\theta_{R})$, are functions of the local covariates $\bm{z}_{k}$, and are modeled using deep neural networks. Therefore, we model the outage state transition for each unit $k$ in (3), i.e., $\mathrm{d}\mathbf{S}_{k}(t)/\mathrm{d}t$, as follows:
@@ III-C Model Estimation
where $\rho>0$ ensures differentiability. More implementation details can be found in Appendix -D.
@@ IV Application: Mobile Generator Deployment
Figure 3: A synthetic example of the mobile generator deployment problem for a system with three cities and five generators ($Q_{w}=5$). The $y$-axis represents the number of outaged households. In this example, the transportation cost is set to $c=1000$, the customer interruption cost is $\tau=50$, and the operational cost is $\gamma=0.2$. Four methods are compared on out‑of‑sample data: the proposed GDF framework (regret = $94.22$), a two‑stage approach (regret = $8992.92$), an online baseline (regret = $790.96$), and the optimal solution with the lowest total cost (regret = $8033.85$). Details of the synthetic data setup are provided in Section V-A.
@@
The mobile generator deployment problem is therefore defined on a directed graph denoted by $(\mathcal{V},\mathcal{E})$, where $\mathcal{V}\coloneqq\mathcal{K}\cup\mathcal{W}$ is the vertex set, and $\mathcal{E}$ represents the edge set defining feasible transportation routes. Each edge $(k,k^{\prime})\in\mathcal{E}$ incurs a transportation cost $c_{kk^{\prime}}$. The decision variables, ${\bm{x}=\{x_{tkk^{\prime}}\}}$, $(k,k^{\prime})\in\mathcal{E}$, $t\in\mathcal{T}$, specifies the transportation schedule, where $x_{tkk^{\prime}}\in\mathbb{Z}_{+}$ represents the number of generators transported from location $k$ to $k^{\prime}$ at time $t$. For simplicity, we assume negligible travel and deployment time, allowing generators to become operational immediately upon arrival to supply power to $N_{g}$ customers.
@@
$q_{tk}=\sum_{t}(\sum_{k^{\prime}\in\mathcal{K}}x_{tk^{\prime}k}-\sum_{k^{\prime}\in\mathcal{K}}x_{tkk^{\prime}}),\quad\forall k\in\mathcal{V}$ (11)
@@
Constraints (9) and (10) establish the initial generator stock at each location. Equation (11) keeps track of the generator stock in each unit. Inequality constraint (12) limits the flow between service units or warehouses in $\mathcal{E}$ to a maximum capacity $C$. Finally, equation (13) ensures flow conservation, requiring that the total inflow of generators equals the total outflow over the planning horizon for each node in the network. As a result, all generators return to the staging areas after the events.
To implement the deployment strategy based on our GDF framework, we first predict outage levels $\hat{Y}_{k}(t)$ for all service units using the neural ODE model specified in (5). The model is learned by minimizing the GDF loss in (6). These global-deision-focused predictions are then integrated into the objective function (8) of the mobile generator deployment problem, and the decisions are derived by solving a mixed-integer linear programming (MILP). Figure 3 illustrates the performance of GDF in a stylized mobile generator deployment problem. As shown, GDF significantly reduces regret compared to two-stage and online baselines. The online algorithm takes a reactive approach, frequently relocating generators and incurring high operational costs, while the two-stage method fails to capture system-wide dynamics, leading to suboptimal decisions. In contrast, GDF’s globally optimized approach ensures more effective and cost-efficient resource allocation, enhancing overall grid resilience with minimal cost.
@@ V-B Experimental Setup
To evaluate the effectiveness of the proposed Global Decision-Focused (GDF) Neural ODE framework, we conduct experiments on two application domains: power line undergrounding and mobile generator deployment as introduced in Section IV. For both problems, predictions for the entire time horizon are made before the occurrence of the hazard, reflecting the practical challenge of collecting data in real time during a developing hazard [53]. Consequently, for both problems, when the hazard surpasses a predefined threshold (1% in our experiments), the ODE model’s initial conditions are set, and decisions are made accordingly.
In the generator deployment problem, we assume a single centralized warehouse as the sole source of generators, with no inter-county transfers. Generators are dispatched from the warehouse to affected counties and must eventually return for refueling, following the framework described in Section V-C1. The online baseline method assumes a one-day lag, making greedy decisions based on data from the previous day, reflecting a realistic delay in data availability for decision-makers without predictive insights. We also report results for different lag values, while all online methods reported in the plots use a fixed lag of one day.
@@
TABLE I: Out-of-Sample Performance for Generator Distribution problem (Synthetic Data). Results are averaged over 3 repeated experiments with standard error (SE) in the brackets.
Model Transportation Cost = 100 Transportation Cost = 500 Transportation Cost = 1000
@@
Ground Truth / 6515.6 (74.4) 0 / 11271 0 / 16271 0
Online (lag = 1) / 8163.7 (293.4) 1648.1 (358.6) / 19630.4 (1514.8) 8358.7 (1576.8) / 33963.7 (3042.2) 17692.1 (3104.4)
Online (lag = 3) / 8978.6 (277.6) 2462.9 (313.8) / 22045.2 (1493.3) 10773.6 (1522.9) / 38378.6 (3020.2) 22106.9 (3050.4)
Online (lag = 5) / 9618.9 (57.3) 4939.9 (111.1) / 22152.2 (498.5) 16497.9 (557.8) / 37818.9 (1074.7) 32164.5 (1135.0)
Two-stage 9336.0 (975.6) 6671.5 (4328.7) 155.8 (33.3) 9336.0 (975.6) 11465.8 (94.6) 194.2 (107.3) 9336.0 (975.6) 16517.8 (133.9) 246.1 (112.7)
GDF 9709.6 (3055.6) 6668.7 (4365.9) 153.0 (19.8) 8671.2 (1553.7) 11428.1 (87.1) 156.4 (44.2) 6981.1 (1874.0) 16395.9 (89.0) 124.2 (36.3)
@@ V-C1 Mobile Generator Deployment Problem
Table I summarizes the out-of-sample performance for the generator deployment problem on synthetic data across three transportation cost factors (100, 500, and 1000). As the transportation cost increases, the improvement in decision quality for GDF compared to the Two-stage approach becomes more apparent. This highlights the importance of scheduling and proactive actions when transportation costs are high, demonstrating the clear advantage of GDF.
To understand how GDF impacts MSE-trained predictive models, see Fig.3. By slightly increasing the predicted outage for city 3, GDF proactively allocated more resources to city 3’s anticipated outage, drawing from the stock allocated to city 1. In contrast, the Two-stage method did not make this adjustment. While this change in prediction is subtle, it results in significantly lower regret. Remarkably, the GDF model achieves an MSE comparable to or better than that of the MSE-trained model in TableI, albeit with higher variance. This shows that improved decision quality was not achieved at the expense of predictive accuracy, but rather through targeted and meaningful adjustments in prediction.
We conducted extensive experiments on synthetic data for the mobile generator deployment problem under various settings, as shown in Fig. 4. The decision quality improvement of GDF relative to the Two-stage method initially increases with transportation cost before eventually declining. This behavior is expected: higher transportation costs highlight the importance of proactive actions. However, when transportation costs become so high that the best strategy is effectively to remain idle, the potential savings for proactive actions diminish. Similarly, we observed that as operational costs and available generators increase, the performance gap between GDF and the other two methods narrows.

### v2

@@ I Introduction
To proactively mitigate power grid disruptions, it is crucial to anticipate the long-term impact of extreme events on power systems and implement preemptive measures accordingly. Accurate forecasting enables utility companies and policymakers to assess risks, prioritize interventions, optimize grid reinforcement, and allocate resources efficiently. By identifying vulnerable regions in advance, decision makers can develop strategies that not only reduce power disruptions but also strengthen overall grid resilience.
@@ I-A Related Work
To overcome these limitations, we propose integrating decision-focused learning (DFL) with predictive modeling. By embedding decision objectives directly into the learning process, our approach aligns predictions with optimization goals, enabling more adaptive and proactive resource allocation and grid reinforcement. This integration enhances grid resilience amid escalating natural hazard risks.
Power Outage Modeling. Accurate power outage forecasting is crucial for enhancing grid reliability and resilience [23, 24]. Various machine learning and statistical methods have been employed to predict outages under different conditions. These approaches incorporate neural networks enriched with environmental factors and semantic analysis of field reports, providing real-time updates and enhancing predictive performance through text analysis.
@@ III-A Neural ODEs for Power Outages
Drawing inspiration from the Susceptible Infectious Recovered (SIR) models commonly used in epidemic modeling [45], we conceptualize power outages within a unit as the spread of a “virus”. In this analogy, outages propagate among customers due to local transmission line failures, while restorations provide lasting resistance to subsequent outages. We formalize this analogy with the following three assumptions: ($i$) The number of unaffected customers $U_{k}(t)$ decreases over time as some customers transition from being unaffected to experiencing outages. This transition is governed by the failure transmission rate, denoted by $\phi(\bm{z}_{k};\theta_{U})$. Inspired by epidemiological transmission rates, this rate quantifies how local conditions – such as weather patterns and other regional factors encapsulated by covariates $\bm{z}_{k}$ – influence the rate at which outages spread within the grid. ($ii$) Conversely, the number of customers with restored power ($R_{k}(t)$) gradually increases as the system operator repairs transmission lines and restores service. This process is captured by the restoration rate, denoted $\phi(\bm{z}_{k};\theta_{R})$. Both transmission and restoration rates, $\phi(\bm{z}_{k};\theta_{U})$ and $\phi(\bm{z}_{k};\theta_{R})$, are functions of the local covariates $\bm{z}_{k}$, and are modeled using deep neural networks. ($iii$) The total number of customers within each unit remains constant throughout the studied period. Based on these assumptions, we model the outage state transition for each unit $k$ in (3), i.e., $\mathrm{d}\mathbf{S}_{k}(t)/\mathrm{d}t$, as follows:
@@ III-C Model Estimation
where $\rho>0$ ensures differentiability. More implementation details can be found in the supplementary material.
@@ IV Application: Mobile Generator Deployment
Figure 3: A synthetic example of the mobile generator deployment problem for a system with three cities and five generators ($Q_{w}=5$). The $y$-axis represents the number of outaged households. In this example, the uniform travel time $\delta_{t}=10$, transportation cost is set to $c=400$, the customer interruption cost is $\tau=1$, and the operational cost is $\gamma=2$. Four methods are compared on out‑of‑sample data: the proposed GDF framework (regret = $183.47$), a two‑stage approach (regret = $488.45$), an online baseline with observation lag of 5 (regret = $2846.13$), and the optimal solution with the groundtruth. Details of the synthetic data setup are provided in Section V-A.
@@
The mobile generator deployment problem is therefore defined on a directed graph denoted by $(\mathcal{V},\mathcal{E})$, where $\mathcal{V}\coloneqq\mathcal{K}\cup\mathcal{W}$ is the vertex set, and $\mathcal{E}$ represents the edge set defining feasible transportation routes. Each edge $(k,k^{\prime})\in\mathcal{E}$ incurs a transportation cost $c_{kk^{\prime}}$, and an associated travel time $\delta_{kk^{\prime}}\in\mathbb{Z}_{+}$, representing the number of time periods required for generators to be transported from location $k$ to location $k^{\prime}$. The decision variables, ${\bm{x}=\{x_{tkk^{\prime}}\}}$, $(k,k^{\prime})\in\mathcal{E}$, $t\in\mathcal{T}$, specifies the transportation schedule, where $x_{tkk^{\prime}}\in\mathbb{Z}_{+}$ represents the number of generators transported from location $k$ to $k^{\prime}$ at time $t$. For simplicity, we assume negligible deployment time, allowing generators to become operational immediately upon arrival to supply power to $N_{g}$ customers.
@@
$q_{tk}=\sum_{k^{\prime}\in\mathcal{V}}\sum_{\begin{subarray}{c}t^{\prime}\in\mathcal{T}\\ t^{\prime}+\delta_{k^{\prime}k}\leq t\end{subarray}}x_{t^{\prime}k^{\prime}k}-\sum_{k^{\prime}\in\mathcal{V}}x_{tkk^{\prime}},\quad\forall k\in\mathcal{V},\,t\in\mathcal{T}$ (11)
@@
Constraints (9) and (10) establish the initial generator stock at each location. Equation (11) tracks the generator stock at each location, accounting for the travel time $\delta$ taken for incoming shipments. Inequality constraint (12) limits the flow between service units or warehouses in $\mathcal{E}$ to a maximum capacity $C$. Finally, equation (13) ensures flow conservation, requiring that the total inflow of generators equals the total outflow over the planning horizon for each node in the network. As a result, all generators return to the staging areas after the events.
To implement the deployment strategy based on our GDF framework, we first predict outage levels $\hat{Y}_{k}(t)$ for all service units using the neural ODE model specified in (5). The model is learned by minimizing the GDF loss defined in (6). These global-deision-focused predictions are then integrated into the objective function (8) of the mobile generator deployment problem, and the decisions are derived by solving a mixed-integer linear programming (MILP). Figure 3 illustrates the optimal transportation schedules derived from the GDF predictions using an MILP solver, compared to the schedules produced by the Two-stage and reactive online approaches in a stylized example of the mobile generator deployment problem.
@@ V-B Experimental Setup
To evaluate the effectiveness of the proposed Global Decision-Focused (GDF) Neural ODE framework, we conduct experiments on two application domains: power line undergrounding [48, 49, 50, 51] and mobile generator deployment as introduced in Section IV. For both problems, predictions for the entire time horizon are made before the occurrence of the hazard, reflecting the practical challenge of collecting data in real time during a developing hazard [52]. Consequently, for both problems, when the hazard surpasses a predefined threshold (1% in our experiments), the ODE model’s initial conditions are set, and decisions are made accordingly.
In the generator deployment problem, we assume a single centralized warehouse as the sole source of generators, with no inter-county transfers. Generators are dispatched from the warehouse to affected counties and must eventually return for refueling, following the setup described in Section V-C1. We also assume a uniform travel time parameter, $\delta_{kk^{\prime}}=\delta_{t},\;\forall(k,k^{\prime})\in\mathcal{E}$, to capture the impact of transportation delays. The online baseline method assumes an observation lag, making greedy observe-then-optimize decisions based on data from previous data, reflecting a realistic delay in data availability for decision makers without predictive insights.
@@
TABLE I: Out-of-Sample Performance for Generator Distribution Problem with Synthetic Data. Results are averaged over 3 repeated experiments with standard error (SE) in brackets.
Model $\delta_{t}$ = 1 $\delta_{t}$ = 5 $\delta_{t}$ = 10
@@
Ground Truth / 10316.9 (73.9) 0 / 11031.7 (69.0) 0 / 12473.7 (63.8) 0
Online (lag = 1) / 63663.2 (3762.5) 53346.3 (3769.9) / 18862.8 (965.7) 7831.1 (916.1) / 19681.8 (2069.3) 7208.1 (2121.0)
Online (lag = 3) / 84769.4 (122.8) 74452.6 (7246.8) / 19816.9 (73.6) 8785.2 (1216.9) / 38378.6 (3020.2) 4962.5 (792.7)
Two-stage 9336.0 (975.6) 10459.8 (90.6) 142.9 (90.3) 9336.0 (975.6) 11506.6 (98.1) 474.9 (95.7) 9336.0 (975.6) 12746.2 (196.7) 272.4 (207.2)
GDF 7204.1 (1804.8) 10409.9 (79.7) 93.0 (11.8) 7531.5 (1805.8) 11452.1 (80.3) 420.3 (24.3) 8907.7 (2645.2) 12601.9 (58.1) 128.2 (49.3)
@@ V-C1 Mobile Generator Deployment Problem
Table I summarizes the out-of-sample performance for the generator deployment problem on synthetic data under different travel time settings ($\delta_{t}=1,5,10$). The proposed method consistently delivers the best decision quality, as reflected in the lowest regret. Notably, as $\delta_{t}$ increases—representing longer transportation delays—the performance improvement of GDF over the two-stage approach becomes increasingly significant. A similar trend is observed under higher transportation cost factors, as shown in Fig. 4, until transportation costs become so high that the best strategy is effectively to remain idle, at which point the improvement from proactive actions with GDF diminishes.
These findings highlight the importance of proactive scheduling and early interventions when facing longer deployment delays or higher transportation costs. It also demonstrates the advantage of GDF in optimizing resource allocation under more challenging, time-sensitive conditions in the face of extreme hazards that threaten grid resilience. As visualized in Fig. 3, the GDF model slightly overestimate outages in critical regions first hit by outage, enabling the reallocation of additional resources to mitigate potential disruptions. While this adjustment in the forecast is subtle, it leads to a significant reduction in regret. Remarkably, the GDF models also achieve MSE comparable to that of the MSE-trained model in Table I, albeit with higher variance. This indicates that improved decision quality was not achieved at the expense of predictive accuracy but rather through targeted and meaningful adjustments in the forecasts.
Additionally, we conducted extensive experiments on synthetic data for the mobile generator deployment problem under various settings, as shown in Fig. 4. We observed that the performance advantage of GDF over the Two-stage and greedy methods narrows as customer interruption costs, operational costs, or the number of available generators increase. This behavior is expected, as higher system resources or reduced flexibility diminish the need for globally optimized interventions—resulting in less improvement with GDF.
@@ VI Conclusion
Our research provides actionable insights for grid resilience practitioners, highlighting the importance of proactive scheduling and early interventions when facing the threat of natural hazards to power systems. It also shows that embedding global-decision-focused learning objectives into predictive models can meaningfully shift forecasting to improve system-wide outcomes. Specifically, by strategically adjusting predicted outages—such as preemptively amplifying forecasts for high-risk or critical regions—the GDF framework enables earlier and more targeted resource allocation decisions. This leads to substantial reductions in overall costs and, consequently, in regret. The benefits are especially pronounced in settings with high transportation costs or delays, where globally optimized forecasts are key for operators to prioritize and allocate limited resources proactively and more effectively across the entire network. In such scenarios, the forecasting model, empowered by decision-aware optimization, acts not just as a passive predictor but as a key lever in optimizing strategic resource deployment, ultimately improving grid resilience and reducing the economic burden of extreme events.